# hide_content

### v2

@@ 5.2 Metrics
Our primary metric is task success rate which is the standard for agent evaluations [12, 32, 6]. VisualWebArena provides an evaluation framework for task success. Task success criteria include achieving an expected final webpage state or receiving a desired response from the language agent. OmniACT does not provide task success rate directly, instead introducing the sequence score and action score metrics. Sequence score evaluates if the output contains the correct high level action (e.g. click or type), but does not check if the action element or parameter (e.g. click [1] or type [parameter]) are correct. Action score evaluates if the output contains both the correct high level action and the correct element or parameter. Thus, for OmniACT we focus our evaluations on the action score as it is more similar to task success rate.
@@ 5.3 Ordering Methods
We apply dimensionality reduction techniques to better capture the spatial relationships. Using t-SNE [21], we reduce the dimensionality with the function $g\colon\langle x,y\rangle\rightarrow z$. The set of values $Z=\{z_{1},z_{2},\ldots,z_{m}\}$ is used to determine the ordering $\sigma_{tsne}$. We use the scikit-learn [14] implementation of t-SNE and keep the default parameters.
@@ 6.2 State-of-the-Art Performance on OmniACT
We achieve a new state-of-the-art performance on Omniact. Due to cost, we look to our previous experiments to pick the best combination of features for our approach. We observe that multimodal representations are still helpful. We find that t-SNE ordering improves performance the best in OmniACT when elements are detected by our model. Koh et al. [12] states that high level actions are easier for a LM to reason with.
@@
  • Element Source: OmniACT obtains elements by searching for icons with Segment Anything [11] and text with EasyOCR. Unfortunately, there are no shared artifacts for their icon detection system. We obtain UI elements through an object detection model and text with EasyOCR[1].
@@
We apply our findings in and achieve more than one-fold increase over the existing best action score. Our results can be found in Table 6.
@@
Table 6:  and x indicates whether the feature is available when building a representation for the model. Ours indicates a high level actions such as click [1]. Code indicates that pyautogui code is directly generated. bold is best and italics are second best. While sequence score only checks for the correct high level action (e.g. click), action score checks for both the correct action and the correct element or parameter (e.g. click [1]). Thus, action score is the most equivalent to task success rate. * is as reported in [9]
@@ 7 Future Work
We focused on the impact of element ordering in a text representation (although element labels re-ordered accordingly in the visual representation as well). The impact of element ordering may or may not generalize to an image only representation. Unfortunately, our results indicated that a visual representation alone was insufficient for web and desktop environments which prevents us from conducting this experiment. However, Yan et al. [27] found that a visual representation in mobile environments was able to achieve comparable performance with and without a text representation. In the future, we hope to experiment with various ordering methods on an image only representation.
@@
In this paper, we explored two benchmarks—VisualWebArena and OmniACT—as web and desktop scenarios. In the future, we hope to explore other benchmarks and settings with our approach. For example,  Xie et al. [26] uses the OS level accessibility tree for desktop agent navigation. We hope to compare our approach against theirs and believe that combining both approaches may lead to further improvements. Additionally, mobile environments often have only pixel-level information [2, 17]; we hope to apply our approach to a mobile benchmark such as Rawles et al. [15].
7.1 Broader Impacts
LM agents have numerous potential benefits: they may improve accessibility for individuals with disabilities, streamline repetitive tasks, and enhance overall productivity in various digital environments. However, there are important concerns to consider regarding privacy, security, and ethical use. Since agents will require access to a user’s desktop or web environment, there is a heightened risk of privacy breaches if data security is not properly managed properly. Additionally, agent automation could be misused, potentially leading to the spread of misinformation or enabling automated cyber-attacks. Ethical issues also emerge around the transparency and accountability of the agent’s actions; these effects will be compounded as agents start to move from research into real-world settings. There is also a potential impact to the labor force as capable LM agents may displace existing human workers. While our work here focuses on evaluating agents in controlled settings, we believe further work is needed to understand the risks of LM agents before they can be deployed responsibly in real-world settings.
@@ A.1 Faster-RCNN Training Details
We trained our model using the detectron2 [25] implementation of faster-rcnn. We did not change much from the default implementation and recognize that there are significant improvements that could be made to the model.
@@ A.3 LM Agent Hyperparameters and Settings
We set our temperature, top-p, and input token limits based on existing works [32, 12, 9]. Prompts for all three backbones contain few-shot examples and use chain-of-thought prompting [22] as described in Koh et al. [12]. In GPT-4v and Llama3 each example is a different message; Gemini-1.5’s context length allowed us to input all examples as a single prompt. We detail our LM Agent hyperparameters in Table 9

### v3

@@ 5.2 Metrics
Our primary metric is task success rate which is the standard for agent evaluations (Koh et al., 2024; Zhou et al., 2023; He et al., 2024). VisualWebArena provides an evaluation framework for task success. Task success criteria include achieving an expected final webpage state or receiving a desired response from the language agent. OmniACT does not provide task success rate directly, instead introducing the sequence score and action score metrics. Sequence score evaluates if the output contains the correct high level action (e.g. click or type), but does not check if the action element or parameter (e.g. click [1] or type [parameter]) are correct. Action score evaluates if the output contains both the correct high level action and the correct element or parameter. Thus, for OmniACT we focus our evaluations on the action score as it is more similar to task success rate.
@@ 5.3 Ordering Methods
We apply dimensionality reduction techniques to better capture the spatial relationships. Using t-SNE (Van der Maaten & Hinton, 2008), we reduce the dimensionality with the function $g\colon\langle x,y\rangle\rightarrow z$. The set of values $Z=\{z_{1},z_{2},\ldots,z_{m}\}$ is used to determine the ordering $\sigma_{tsne}$. We use the scikit-learn (Pedregosa et al., 2011) implementation of t-SNE and keep the default parameters.
@@ 6.2 State-of-the-Art Performance on OmniACT
We achieve a new state-of-the-art performance on Omniact. Due to cost, we look to our previous experiments to pick the best combination of features for our approach. We observe that multimodal representations are still helpful. We find that t-SNE ordering improves performance the best in OmniACT when elements are detected by our model. Koh et al. (2024) states that high level actions are easier for a LM to reason with.
@@
  • Element Source: OmniACT obtains elements by searching for icons with Segment Anything (Kirillov et al., 2023) and text with EasyOCR. Unfortunately, there are no shared artifacts for their icon detection system. We obtain UI elements through an object detection model and text with EasyOCR (AI, 2020).
@@
We apply our findings and achieve more than one-fold increase over the existing best action score. Our results can be found in Table 6.
@@
Table 6:  and x indicates whether the feature is available when building a representation for the model. Ours indicates a high level actions such as click [1]. Code indicates that pyautogui code is directly generated. bold is best and italics are second best. While sequence score only checks for the correct high level action (e.g. click), action score checks for both the correct action and the correct element or parameter (e.g. click [1]). Thus, action score is the most equivalent to task success rate. * is as reported in Kapoor et al. (2024)
@@ 7 Future Work
We focused on the impact of element ordering in a text representation (although element labels re-ordered accordingly in the visual representation as well). The impact of element ordering may or may not generalize to an image only representation. Unfortunately, our results indicated that a visual representation alone was insufficient for web and desktop environments which prevents us from conducting this experiment. However, Yan et al. (2023) found that a visual representation in mobile environments was able to achieve comparable performance with and without a text representation. In the future, we hope to experiment with various ordering methods on an image only representation.
@@
In this paper, we explored two benchmarks—VisualWebArena and OmniACT—as web and desktop scenarios. In the future, we hope to explore other benchmarks and settings with our approach. For example,  Xie et al. (2024) uses the OS level accessibility tree for desktop agent navigation. We hope to compare our approach against theirs and believe that combining both approaches may lead to further improvements. Additionally, mobile environments often have only pixel-level information (Chen et al., 2020; Ross et al., 2020); we hope to apply our approach to a mobile benchmark such as Rawles et al. (2023).
@@ A.1 Faster-RCNN Training Details
We trained our model using the detectron2 Wu et al. (2019) implementation of faster-rcnn. We did not change much from the default implementation and recognize that there are significant improvements that could be made to the model.
@@ A.3 LM Agent Hyperparameters and Settings
We set our temperature, top-p, and input token limits based on existing works Zhou et al. (2023); Koh et al. (2024); Kapoor et al. (2024). Prompts for all three backbones contain few-shot examples and use chain-of-thought prompting Wei et al. (2023) as described in Koh et al. (2024). In GPT-4v and Llama3 each example is a different message; Gemini-1.5’s context length allowed us to input all examples as a single prompt. We detail our LM Agent hyperparameters in Table 9